# Triplet Interaction Improves Graph Transformers: Accurate Molecular Graph Learning with Triplet Graph Transformers

### v1

@@ 4 Experiments
Our experiments are designed to validate several key aspects of our proposed model and training approach. Firstly, we demonstrate the performance and scalability of our approach on large quantum chemistry datasets PCQM4Mv2 (Hu et al., 2021) and OC20 (Chanussot et al., 2021). Next, we evaluate the transfer learning capabilities of our models, by finetuning our task predictor from PCQM4Mv2 to related quantum chemistry tasks on the QM9 (Ramakrishnan et al., 2014) dataset. We also transfer our distance predictor from PCQM4Mv2 to molecular property prediction on MOLPCBA (Hu et al., 2020) and drug discovery dataset LIT-PCBA (Tran-Nguyen et al., 2020). Finally, we showcase the utility of our triplet interaction mechanisms for graph learning in general by evaluating it on the traveling salesman problem task on the TSP dataset by Dwivedi et al. (2020). The PyTorch (Paszke et al., 2019) library was used to implement the models. The training was done with mixed-precision computation on 4 nodes, each with 8 NVIDIA Tesla V100 GPUs (32GB RAM/GPU), and two 20-core 2.5GHz Intel Xeon CPUs (768GB RAM). The hyperparameters, training, and dataset details are provided in the appendix. Our code is available at https://github.com/shamim-hussain/tgt.
@@ 4.1 Large-scale Quantum Chemical Prediction
PCQM4Mv2 PCQM4Mv2 comprising 4 million molecules, is a part of the OGB-LSC datasets (Hu et al., 2021). The primary objective is predicting the the HOMO-LUMO gap. The performance of the distance predictor is tuned on a random subset of 5% of the training data which we call validation-3d. The training of our TGT-At model takes A100 GPU-days, slightly less than the training time of UniMol+ (Lu et al., 2023), which takes 40 A100 GPU-days.
The results of our experiments are presented in Table 4 in terms of Mean Absolute Error (MAE) in meV unit. Our TGT-At model achieves the best performance on the PCQM4Mv2 dataset, outperforming the previous SOTA UniMol+ model by a significant margin of 2.2 meV. It’s worth highlighting that UniMol+ uses RDKit coordinates as input which is optional for our model. We see that even without RDKit coordinates, i.e., with a pure neural approach, our model outperforms UniMol+ by a fair margin. Hence, we currently hold the top positions on the PCQM4Mv2 leaderboard for both RDKit-aided and pure neural approaches. The TGT-Ag model also exhibits good performance, securing the second-best position after TGT-At. TGT-Agx2 reduces parameter count by half by sharing parameters in consecutive layers yet still outperforms UniMol+. Our performance gains over other models stem from two factors – not only a superior architecture but also better training and inference. This is evidenced by the success of a basic EGT 2-stage model under our training and inference paradigm.
@@
Uncertainty Estimation Our stochastic inference method allows us to draw multiple sample predictions for each input, which can be used to estimate the uncertainty of our predictions, e.g., by looking at the spread, and using the reciprocal of standard deviation as a confidence measure, as shown in Figure 4. Here we normalize the confidence to the range for better visualization, and we plot the MAE and EwT of validation graphs, filtered by the confidence threshold. We can see that performance monotonically improves (lower MAE and higher EwT) with higher confidence. This shows that the confidence of a prediction has a strong positive correlation with its accuracy. This can be useful for real-world applications like drug discovery and material design.
@@ 4.2 Transfer Learning
Finetuning on QM9 To highlight the transfer of knowledge to related quantum chemical prediction tasks we finetune the task predictor from PCQM4Mv2 on the QM9 (Ramakrishnan et al., 2014) dataset. The ground truth 3D coordinates are provided on this dataset which can be used during inference, so the distance predictor is not required. We report finetuning performance on a subset of 7 tasks from the 12 tasks in QM9. We get comparable results with TGT-Ag and TGT-At, so we only report results for TGT-Ag to save compute. We see that TGT-Ag archives SOTA results and outperforms other models by a significant margin in predicting the HOMO (), LUMO(), and the HOMO-LUMO gap (). This is because these tasks are directly related to the pretraining task. We also achieve SOTA results for and and perform satisfactorily on the other two tasks – demonstrating a positive transfer of knowledge to these tasks. Notably, we outperform Transformer-M, another transformer model pretrained on PCQM4Mv2.
@@
Baseline Methods: GraphMVP (Liu et al., 2021a), GEM (Fang et al., 2021), 3D Infomax (Stärk et al., 2022), 3D-MGP (Jiao et al., 2022), Schnet (Schütt et al., 2017), PhysNet (Unke & Meuwly, 2019), Cormorant (Anderson et al., 2019), DimeNet++ (Gasteiger et al., 2020a), PaiNN (Schütt et al., 2021), EGNN (Satorras et al., 2021), NoisyNode (Godwin et al., 2021), SphereNet (Liu et al., 2021b), SEGNN (Brandstetter et al., 2021), EQGAT (Le et al., 2022), SE(3)-Transformer (Fuchs et al., 2020), TorchMD-Net (Thölke & De Fabritiis, 2022), Equiformer (Thölke & De Fabritiis, 2021), Transformer-M (Luo et al., 2022)
@@
| GIN-VNpretrain | 3.4M | 29.02(0.17) |
| Graphormer-FLAGpretrain | 119.5M | 31.40(0.34) |
| EGTpretrain | 110.8M | 29.61(0.24) |
@@
Baseline Models: DeeperGCN(Li et al., 2020), PNA (Corso et al., 2020), DGN (Beani et al., 2021), GINE(Brossard et al., 2020), PHC-GNN (Le et al., 2021), GIN-VN (Xu et al., 2018), Graphormer (Ying et al., 2021), EGT (Hussain et al., 2022), Virtual Nodes (-VN) (Gilmer et al., 2017), FLAG (Kong et al., 2020)
| Avg. Test | |
| Model | ROC-AUC(%) |
| GCN | 72.3 |
| GAT | 75.2 |
| FP-GNN | 75.9 |
| EGT | 66.7 |
| EGTpretrain | 78.9 |
| GEM | 76.6 |
| GEMpretrain | 78.4 |
| GEM-2 | 77.6 |
| GEM-2pretrain | 81.5 |
| EGT+RDKit | 81.2 |
| EGT+TGT-At-DP | 81.5 |
Baseline Models: GCN (Kipf & Welling, 2016), GAT (Veličković et al., 2017) and FP-GNN (Cai et al., 2022a), EGT (Hussain et al., 2022), GEM (Fang et al., 2021), GEM-2 (Liu et al., 2022a)
| #Param | #Param | |
| Model | 100K | 500K |
@@
Molecular Property Prediction Since 3D geometric information is valuable for molecular property prediction, we use our pretrained distance predictor (without RDKit) to provide estimations of interatomic distances to models on the MOLPCBA (Hu et al., 2020) molecular property prediction and the LIT-PCBA (Tran-Nguyen et al., 2020) drug discovery benchmarks. These datasets do not have ground truth 3D information. So, we do not finetune the distance predictor on them but rather use it as a frozen feature extractor.
The results for MOLPCBA are presented in Table 8 in terms of test Average Precision (%) which evaluates the performance of the model in a multitask setting of predicting 128 different binary molecular properties. We see that both EGT and TGT-Ag models trained from scratch with RDKit coordinates get good results but if we use our pretrained TGT-AT-DP (“-DP” stands for distance predictor), we get the best results. We also see that our model outperforms the SOTA pretrained Graphormer model by a significant margin. On the LIT-PCBA dataset, we report on an average ROC-AUC (%) on 7 separate tasks of predicting interactions with proteins (a breakdown is provided in the appendix). We see that EGT with TGT-At-DP matches the SOTA pretrained GEM-2. Both of these experiments demonstrate that our pretrained TGT-At-DP can provide valuable 3D information to the task predictor, even though it is trained on a different dataset. We also see that our TGT-At-DP which is trained on DFT coordinates can provide more useful 3D information than RDKit coordinates.
@@ 4.3 Traveling Salesman Problem
We also show the utility of our proposed pair-to-pair communication mechanism beyond molecular graphs and for general-purpose graph learning by targetting the Traveling Salesman Problem benchmark dataset by Dwivedi et al. (2020) which consists of 12000 K-NN graphs of 50-500 2D points in the unit square. The task is to predict which of the edges of the K-NN graph are part of the optimal tour. According to the specification of the dataset, the task must be performed with a given parameter budget of either 100K or 500K. The results are presented in Table 8 in terms of test F1 score (%) where We get a significant improvement when we use our TGT-Ag and TGT-At models. This shows that our pair-to-pair communication mechanism is very useful for solving the TSP task. We do not evaluate the TGT-At model for the larger 500K parameter budget due to memory constraints. We show the performance of the TGT-Ag model can be further improved by using repeated layers with shared parameters, dubbed as TGT-Agx2, TGT-Agx3, and TGT-Agx4. This shows the effectiveness of our triplet interaction module in an iterative setting.
@@ 4.4 Ablation Study
Table 10 compares triplet interaction with axial attention and triangular update, which can be thought of as ablated variants; we also compare with ungated variants. We compare the cross-entropy loss of the distance predictor with different mechanisms on PCQM4Mv2 validation-3D set. We normalize training time with respect to no triplet scenario. We see that triplet attention performs best but is expensive. Triplet aggregation too performs better than both axial attention and triangular update, yet is more efficient. The ungated variants perform slightly worse but are also slightly more efficient. We also see that the model without triplet interaction performs the worst by a significant margin, which shows its importance for distance prediction.
@@
| Triplet | Att. | Update | Triplet | Agg. | Triplet | Att. | |
| Inter. | Agg. | Att. | |||||
| Cross-Ent.() | 1.270 | 1.231 | 1.225 | 1.226 | 1.218 | 1.207 | 1.199 |
| Time/Epoch() | 1.0 | 2.2 | 1.8 | 1.6 | 1.7 | 3.1 | 3.3 |
@@
Table 10 shows a detailed ablation study to test the effectiveness of our triplet interaction and other proposed optimizations within TGT. Results are shown for a smaller 12-layer model on PCQM4Mv2. We see that to take full advantage of the input 3D information (RDKit coordinates), we also need the denoising distance prediction objective. Local smoothing of the input noise improves this process, and source dropout performs better than attention dropout. Incorporating DFT distance predictor, and pretraining on noisy DFT coordinates lead to significant improvements. Finally, a significant leap comes from triplet attention.
@@ 5 Conclusion
In this work, we introduce the Triplet Graph Transformer (TGT) architecture, which enables direct communication between neighboring pairs in a graph. This significantly improves the modeling of geometric dependencies in molecular graphs. We proposed two forms of triplet interactions – an attention-based mechanism with maximum expressivity, and an aggregation-based mechanism with greater efficiency and scalability. Additionally, we put forth a two-stage framework involving separate distance predictor and property predictor models. Our distance predictor directly predicts interatomic distances from 2D graphs during inference, eliminating the need for property prediction on 2D information only. The three-stage training methodology with a stochastic inference scheme enables fast and accurate predictions, significantly advancing over previous iterative refinement approaches, and allows for uncertainty quantification in the prediction. Through extensive experiments, we demonstrate state-of-the-art predictive accuracy on quantum chemical datasets and the transferability of our distance predictor to molecular property prediction. Moreover, the superior performance of TGT on the TSP task indicates the broad applicability of our proposed triplet interactions. In future work, we plan to explore the use of our triplet interaction mechanism for other graph learning tasks. We also plan to evaluate its performance in other molecular graph learning tasks like molecule and conformation generation. Additionally, we aim to further investigate improving the compute and memory efficiency of triple interaction.
6 Broader Impact
Our work aims to improve chemical property prediction, which can accelerate the discovery of new beneficial materials and drugs. This has the potential to greatly benefit society by enabling the development of cheaper, safer, and more effective medicines as well as advanced materials for clean energy and other applications. Our model can be used to efficiently solve the Traveling Salesman Problem for improved routing and logistics planning in transportation. However, though increased automation can reduce costs, over-reliance on AI systems without human oversight raises concerns about accountability and bias. We plan to make the code open-source so that domain experts can further improve and validate the models before real-world deployment.
@@ Appendix A Additional Details about Related Works
Graph Transformers Before pure graph transformers, the self-attention mechanism was used to boost the expressivity of localized message-passing Graph Neural Networks (GNNs) - for example, GraphTrans (Wu et al., 2021) and GSA (Wang & Deng, 2021) used global self-attention to improve long-range information exchange in GNNs. GROVER (Rong et al., 2020) utilized GNNs to generate query, key, and value matrices for self-attention, enabling pretraining on molecular graphs. Graphormer (Ying et al., 2021) was the first to propose a pure graph transformer with graph-specific positional encoding and showed superior performance on molecular property prediction tasks. EGT (Hussain et al., 2022) extended the transformer framework to include pairwise/edge channels and proposed a general framework for graph learning including edge-related and pairwise tasks. GEM-2 (Liu et al., 2022a) extended the notion of pairwise channels to include higher-order channels to account for many body interactions in molecular graphs. GRPE (Park et al., 2022) proposed a more expressive relative positional encoding scheme. TokenGT (Kim et al., 2022) proposed to include both nodes and edges as tokens in the transformer. UniMol and UniMol+ (Lu et al., 2023) use transformer backbones with pairwise channels, similar to EGT, for molecular property prediction. GPS (Rampášek et al., 2022) proposed a framework to combine message-passing and self-attention mechanisms and GPS++ (Masters et al., 2022) tuned these choices to achieve SOTA performance on PCQM4Mv2. Our TGT model is a pure transformer architecture for graph learning with a novel triplet interaction mechanism for direct edge-to-edge communication which is computationally much cheaper than higher-order channels used in GEM-2 while still being expressive enough to capture geometric information required for molecular property prediction.
@@
Appendix B Pair to Pair Communication in Previous Works

### v2

@@ 4 Experiments
Our experiments are designed to validate several key aspects of our proposed model and training approach. Firstly, we demonstrate the performance and scalability of our approach on large quantum chemistry datasets PCQM4Mv2 (Hu et al., 2021) and OC20 (Chanussot et al., 2021). Next, we evaluate the transfer learning capabilities of our models, by finetuning our task predictor from PCQM4Mv2 to related quantum chemistry tasks on the QM9 (Ramakrishnan et al., 2014) dataset. We also transfer our distance predictor from PCQM4Mv2 to molecular property prediction on MOLPCBA (Hu et al., 2020) and drug discovery dataset LIT-PCBA (Tran-Nguyen et al., 2020). Finally, we showcase the utility of our triplet interaction mechanisms for graph learning in general by evaluating it on the traveling salesman problem task on the TSP dataset by Dwivedi et al. (2020). The PyTorch (Paszke et al., 2019) library was used to implement the models. The training was done with mixed-precision computation on 4 nodes, each with 8 NVIDIA Tesla V100 GPUs (32GB RAM/GPU), and two 20-core 2.5GHz Intel Xeon CPUs (768GB RAM). The hyperparameters, training, and dataset details are provided in Appendix E. Our code is available at https://github.com/shamim-hussain/tgt.
@@ 4.1 Large-scale Quantum Chemical Prediction
PCQM4Mv2 PCQM4Mv2 comprising 4 million molecules, is a part of the OGB-LSC datasets (Hu et al., 2021). The primary objective is predicting the HOMO-LUMO gap. The performance of the distance predictor is tuned on a random subset of 5% of the training data which we call validation-3D. The training of our TGT-At model takes A100 GPU-days, slightly less than the training time of UniMol+ (Lu et al., 2023), which takes 40 A100 GPU-days.
The results of our experiments are presented in Table 2 in terms of Mean Absolute Error (MAE) in meV unit. Our TGT-At model achieves the best performance on the PCQM4Mv2 dataset, outperforming the previous SOTA UniMol+ model by a significant margin of 2.2 meV. It is worth highlighting that UniMol+ uses RDKit coordinates as input which is optional for our model. We see that even without RDKit coordinates, i.e., with a pure neural approach, our model outperforms UniMol+ by a fair margin. Hence, we currently hold the top positions on the PCQM4Mv2 leaderboard for both RDKit-aided and pure neural approaches. The TGT-Ag model also exhibits good performance, securing the second-best position after TGT-At. TGT-Agx2 reduces parameter count by half by sharing parameters in consecutive layers yet still outperforms UniMol+. Our performance gains over other models stem from two factors – not only a superior architecture but also better training and inference. This is evidenced by the success of a basic EGT 2-stage model under our training and inference paradigm.
@@
Uncertainty Estimation Our stochastic inference method allows us to draw multiple sample predictions for each input, which can be used to estimate the uncertainty of our predictions by looking at the spread of the samples. Specifically, we use the reciprocal of standard deviation as a confidence measure, as shown in Figure 4. For better visualization, we normalize the confidence measure to the range . We plot the MAE and EwT of validation graphs, filtered by confidence threshold. We can see that performance monotonically improves with higher confidence – evidenced by lower MAE and higher EwT. This shows that the confidence of a prediction has a strong positive correlation with its accuracy. This can be useful for real-world applications like drug discovery and material design.
@@ 4.2 Transfer Learning
Finetuning on QM9 To highlight the transfer of knowledge to related quantum chemical prediction tasks we finetune the task predictor from PCQM4Mv2 on the QM9 (Ramakrishnan et al., 2014) dataset. The ground truth 3D coordinates are provided on this dataset which can be used during inference, so the distance predictor is not required. We report finetuning performance on a subset of 7 tasks from the 12 tasks in QM9 in Table 5. We get comparable results with TGT-Ag and TGT-At, so we only report results for TGT-Ag to save compute. We see that TGT-Ag archives SOTA results and outperforms other models by a significant margin in predicting the HOMO (), LUMO(), and the HOMO-LUMO gap (). This is because these tasks are directly related to the pretraining task. We also achieve SOTA results for and and perform satisfactorily on the other two tasks – demonstrating a positive transfer of knowledge to these tasks. Notably, we outperform Transformer-M (Luo et al., 2022), another transformer model pretrained on PCQM4Mv2 due to our novel triplet interaction mechanism.
@@
Liu et al. (2021a), Fang et al. (2021), Stärk et al. (2022), Jiao et al. (2022), Unke & Meuwly (2019), Schütt et al. (2017), Anderson et al. (2019), Gasteiger et al. (2020a), Schütt et al. (2021), Satorras et al. (2021), Godwin et al. (2021), Liu et al. (2021b), Brandstetter et al. (2021), Le et al. (2022), Fuchs et al. (2020), Thölke & De Fabritiis (2022), Thölke & De Fabritiis (2021), Luo et al. (2022)
@@
| GIN-VN | 3.4M | 29.02(0.17) |
| Graphormer-FLAG | 119.5M | 31.40(0.34) |
| EGT | 110.8M | 29.61(0.24) |
@@
| Test F1(%) | Test F1(%) | |
| Model | (#Param100K) | (#Param500K) |
@@
Molecular Property Prediction Since 3D geometric information is valuable for molecular property prediction, we use our pretrained distance predictor (without RDKit) to provide estimations of interatomic distances to models on the MOLPCBA (Hu et al., 2020) molecular property prediction and the LIT-PCBA (Tran-Nguyen et al., 2020) drug discovery benchmarks. These datasets do not have ground truth 3D information. So, we do not finetune the distance predictor on them but rather use it as a frozen feature extractor. The task predictor is trained from scratch and takes the predicted distances as input.
The results for MOLPCBA are presented in Table 6 in terms of test Average Precision (%) which evaluates the performance of the model in a multitask setting of predicting 128 different binary molecular properties. We see that both EGT and TGT-Ag models trained from scratch with RDKit coordinates get good results but if we use our pretrained TGT-AT-DP (“-DP” stands for distance predictor), we get the best results. We also see that our model outperforms the SOTA pretrained Graphormer model by a significant margin.
On the LIT-PCBA dataset, we report on an average ROC-AUC (%) on 7 separate tasks of predicting interactions with proteins in Table 7 (a breakdown is provided in Table 14 in the appendix). We see that EGT with TGT-At-DP matches the SOTA pretrained GEM-2. Both of these experiments demonstrate that our pretrained TGT-At-DP can provide valuable 3D information to the task predictor, even though it is trained on a different dataset. We also see that our TGT-At-DP which is trained on DFT coordinates can provide more useful 3D information than RDKit coordinates.
@@ 4.3 Traveling Salesman Problem
We also show the utility of our proposed triplet interaction mechanism beyond molecular graphs and for general-purpose graph learning by targetting the Traveling Salesman Problem benchmark dataset by Dwivedi et al. (2020) which consists of 12000 K-NN graphs of 50-500 2D points in the unit square. The task is to predict which of the edges of the K-NN graph are part of the optimal tour. Since distance prediction is not required, we train a single-stage model that performs binary edge classification, and no pretraining/fine-tuning is involved. The model receives both coordinates (node features) and pairwise distances (edge features) as input. According to the specification of this benchmarking dataset, the task must be performed with a given parameter budget of either 100K or 500K.
The results are presented in Table 8 in terms of test F1 score (%) where We get a significant improvement when we use our TGT-Ag and TGT-At models. This shows that our triplet interaction mechanism is very useful for solving the TSP task. We do not evaluate the TGT-At model for the larger 500K parameter budget due to memory constraints. We show the performance of the TGT-Ag model can be further improved by using repeated layers with shared parameters, dubbed as TGT-Agx2, TGT-Agx3, and TGT-Agx4. This shows the effectiveness of our triplet interaction module in an iterative setting.
@@ 4.4 Ablation Study
In Table 9, we compare our proposed triplet interaction methods with the previously proposed 3rd-order mechanisms – axial attention and triangular update, which can be thought of as ablated variants of our method. We also compare with ungated variants of our methods (i.e., without gating terms). We compare the cross-entropy losses of the distance predictors (a good indicator of the downstream performance) with different 3rd-order interaction mechanisms on the PCQM4Mv2 validation-3D set. We normalize training time with respect to the no 3rd-order interaction scenario. We see that triplet attention performs best but is expensive. Triplet aggregation performs better than both axial attention and triangular update, yet is more efficient. The ungated variants perform slightly worse but are also slightly more efficient. We also see that the model with no 3rd-order interaction performs the worst by a significant margin, which shows its importance for distance prediction.
@@
| 3rd- | Att. | Update | Triplet | Agg. | Triplet | Att. | |
| Order | Agg. | Att. | |||||
| Cross-Ent. | 1.270 | 1.231 | 1.225 | 1.226 | 1.218 | 1.207 | 1.199 |
| Time/Epoch | 1.0 | 2.2 | 1.8 | 1.6 | 1.7 | 3.1 | 3.3 |
@@
Table 10 shows a detailed ablation study to test the effectiveness of our triplet interaction mechanism and other proposed optimizations within our model and training framework. Results are shown for a smaller 12-layer model on PCQM4Mv2. We see that to take full advantage of the input 3D information (e.g., RDKit coordinates), we also need the denoising distance prediction objective. Local smoothing of the input noise improves this process. Source dropout proves to be a better alternative to attention dropout. Incorporating DFT distance predictor, and pretraining on noisy DFT coordinates both lead to significant improvements individually and even greater improvements when combined. Finally, a significant leap comes from triplet attention.
5 Limitations
The main limitations of TGT lie in the computational complexity () of the triplet interaction mechanism which is higher than the complexity of base transformers like EGT. This is because, while EGT only considers pairwise interactions, TGT considers 3rd order interactions. However, we predict that this disadvantage can be alleviated by incorporating sparse and/or low-rank interactions. In this work, we focus on maximizing the performance of the model, and any exploration of the trade-off between performance and complexity is left for future work.
@@ 6 Conclusion
In this work, we introduce the Triplet Graph Transformer (TGT) architecture, which incorporates the 3rd-order triplet interaction mechanism. This significantly improves the modeling of geometric dependencies in graph transformers. We proposed two forms of triplet interactions – an attention-based mechanism with maximum expressivity, and an aggregation-based mechanism with greater efficiency and scalability. Additionally, we put forth a two-stage framework involving separate distance predictor and property predictor models for molecular graphs. Our distance predictor directly predicts interatomic distances from 2D graphs during inference, eliminating the need for property prediction on 2D information only. The three-stage training methodology with a stochastic inference scheme enables fast and accurate predictions, significantly advancing over previous iterative refinement approaches, and allows for uncertainty quantification in the prediction. Through extensive experiments, we demonstrate state-of-the-art predictive accuracy on quantum chemical datasets and the transferability of our distance predictor to molecular property prediction. Moreover, the superior performance of TGT on the TSP task indicates the broad applicability of our proposed triplet interactions. In future work, we plan to explore the use of our triplet interaction mechanism for other graph learning tasks. We also plan to evaluate its performance in other molecular graph learning tasks like molecule and conformation generation. Additionally, we aim to further investigate improving the compute and memory efficiency of triple interaction.
Acknowledgements
This work was supported by the Rensselaer-IBM AI Research Collaboration, part of the IBM AI Horizons Network.
Impact Statement
Our work aims to improve chemical property prediction, which can accelerate the discovery of new beneficial materials and drugs. This has the potential to greatly benefit society by enabling the development of cheaper, safer, and more effective medicines as well as advanced materials for clean energy and other applications. Our model can be used to efficiently solve the Traveling Salesman Problem for improved routing and logistics planning in transportation. However, though increased automation can reduce costs, over-reliance on AI systems without human oversight raises concerns about accountability and bias. We have made the code open-source so that domain experts can further improve and validate the models before real-world deployment.
@@ Appendix A Additional Details about Related Works
Graph Transformers Before pure graph transformers, the self-attention mechanism was used to boost the expressivity of localized message-passing Graph Neural Networks (GNNs) - for example, GraphTrans (Wu et al., 2021) and GSA (Wang & Deng, 2021) used global self-attention to improve long-range information exchange in GNNs. GROVER (Rong et al., 2020) utilized GNNs to generate query, key, and value matrices for self-attention, enabling pretraining on molecular graphs. These hybrid approaches were followed by a new research interest in pure graph transformers. SAN (Kreuzer et al., 2021) utilized Laplace Positional Encodings (LPE) in a global self-attention based graph transformer. Graphormer (Ying et al., 2021) proposed graph-specific relative positional encodings and showed superior performance on molecular property prediction tasks. EGT (Hussain et al., 2022) extended the transformer framework to include pairwise/edge channels and proposed a general framework for graph learning including edge-related and pairwise tasks. GEM-2 (Liu et al., 2022a) extended the notion of pairwise channels to include higher-order channels to account for many body interactions in molecular graphs. GRPE (Park et al., 2022) proposed a more expressive relative positional encoding scheme for graphs. TokenGT (Kim et al., 2022) proposed to include both nodes and edges as tokens in the transformer. UniMol and UniMol+ (Lu et al., 2023) use transformer backbones with pairwise channels, similar to EGT, for molecular property prediction. GPS (Rampášek et al., 2022) proposed a framework to combine message-passing and self-attention mechanisms and GPS++ (Masters et al., 2022) tuned these choices to achieve SOTA performance on PCQM4Mv2. Our TGT model is a pure transformer architecture for graph learning with novel triplet interaction mechanisms for 3rd-order interaction between neighboring pairs which is computationally much cheaper than higher-order channels used in GEM-2 while still being expressive enough to capture geometric information required for molecular property prediction.
@@
Appendix B Third-order Interactions in Previous Works